OpenStudy (anonymous):

i was think -7 but I think i'm wrong....Line AB contains points A (8, -4) and B (1, -5). The slope of line AB is -7 -1/7 7 1/7

OpenStudy (fanduekisses):

Slope =\[\frac{ y_{2} - y_{1} }{ x_{2}-x_{1}}\]

OpenStudy (fanduekisses):

so\[\frac{ -5-(-4) }{ 1-8 }= \frac{ -1 }{ -7 }\]

OpenStudy (fanduekisses):

which equals \[\frac{ 1 }{ 7 }\]

OpenStudy (fanduekisses):

because a negative divided by a negative equals positive

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=-5-(-4)/1-8) m=-1/-7 m=1/7

OpenStudy (anonymous):

Thank you!!!!!!!!!!!